### v1

# Three More Generalized Friendship Paradoxes

Abstract
Generalized friendship paradoxes occur when, on average, our friends have more of some attribute than us. These paradoxes are relevant to many aspects of human interaction, notably in social science and epidemiology. Here, we derive new theoretical results concerning the inevitability of a paradox arising, using a linear algebra perspective. Following the seminal 1991 work of Scott L. Feld, we consider two distinct ways to measure and compare averages, which may be regarded as local and global. For local averaging, we show that a generalized friendship paradox holds for nonbacktracking eigenvector centrality. In this sense, our friends are always more important as us, on average. For global averaging, we define loneliness as the reciprocal of the number of friends and show that for this attribute the generalized friendship paradox always holds in reverse. In this sense, we are always more lonely, on average, than our friends. We also derive a global averaging paradox result for the case where the arithmetic mean is replaced by the geometric mean.
1 Motivation
In 1991, the sociologist Scott L. Feld pointed out a sampling bias phenomenon that applies to any unweighted, undirected network [13]. In the context of social networks, on average, our friends have more friends than we do. This effect has subsequently become known as the Friendship Paradox. With regard to our online interactions, it has been argued that the effect distorts our perceptions [22] and contributes to our dissatisfaction [4]. On a more positive note, the paradox may be exploited in the design of strategies to monitor or control the spread of information or disease [3, 9, 8, 14, 16, 28, 29, 36], including vaccination programmes. More generally, it may be used to sample more efficiently from the tail of a degree distribution [34].
In [11] Eom and Jo looked at a Generalized Friendship Paradox: given an externally derived attribute, on average, do our neighbours have more of this attribute than we do? For example, in a scientific collaboration network, do our coauthors have a higher citation count than us on average? Such issues have subsequently been studied in a range of empirical tests, leading to concepts such as the activity and virality paradoxes [21], the H-index paradox [2], the happiness paradox [36] and the enmity paradox [15]. A further direction of research focuses attention on specific classes of networks, as in [7, 12, 23, 24, 37].
The original work of Feld [13] set up two distinct types of averaging process. Following [15] we will distinguish between them by using the terms global and local. (We note that the terms “ego” and “alter” are used in [28].) Among the empirical tests for paradox effects across various application domains, some emphasize local averaging [2, 10, 4, 7, 26, 32, 40, 41] and others global averaging [18, 21, 33]. In this work we contribute to the understanding of both types of generalized friendship paradox by highlighting new cases where a paradox is guaranteed to hold for any network. We thereby add to the theoretical inevitability results on global averaging in [13, 20] and on local averaging in [7, 11, 19]. We provide three main contributions:
  • Theorem 3.1 shows that a local averaging paradox always holds for nonbacktracking eigenvector centrality.
  • Theorem 5.1 shows that a global averaging paradox always holds for loneliness (defined as the reciprocal of degree),
  • Theorem 6.1 shows that a global geometric mean paradox always holds for degree.
In deriving results, to minimize confusion we will use the phrase “inequality” for expressions that arise when we compare related quantities and reserve the phrase “paradox” for a result showing that the inequality holds for all networks.
2 Definitions and Existing Results
We consider undirected, unweighted, connected networks with $n$ nodes and no self-loops. We let $A\in\mathbb{R}^{n\times n}$ denote the adjacency matrix. So $A$ is a symmetric binary matrix with zero diagonal such that $a_{ij}=1$ if nodes $i$ and $j$ are connected and $a_{ij}=0$ otherwise.
We define the degree vector $\mathbf{d}\in\mathbb{R}^{n}$ by
$\mathbf{d}=A\mathbf{1},$
where $\mathbf{1}\in\mathbb{R}^{n}$ is the vector of ones. So $d_{i}$ is the degree of node $i$. In a social network context, $d_{i}$ measures the number of friends that nodes $i$ possesses. We use $\mathbf{d}^{-1}$ to denote the vector with $i$th element $1/d_{i}$; so $\mathbf{d}^{-1}$ contains the reciprocal degrees. We use $\mathbf{d}^{-T}$ to denote the transpose of $\mathbf{d}^{-1}$. Hence, when $\mathbf{x}\in\mathbb{R}^{n}$ we have $\mathbf{d}^{-T}\mathbf{x}=\sum_{i=1}^{n}x_{i}/d_{i}$.
2.1 Global Averages
By definition, the (arithmetic) average degree, or average number of friends, taken over all nodes is given by $\mathbf{1}^{T}\mathbf{d}/n$. Suppose instead that we consider all nodes and compute the average number of friends-of-friends. In this computation, every node appears as a friend $d_{i}$ times, and each time contributes its $d_{i}$ friends to the sum, giving a total of $\mathbf{d}^{T}\mathbf{d}$. The number of terms in this sum is $\mathbf{1}^{T}\mathbf{d}$, since every edge contributes twice. It follows that the difference between the friend-of-friend average and the friend average is
$\frac{\mathbf{d}^{T}\mathbf{d}}{\mathbf{1}^{T}\mathbf{d}}-\frac{\mathbf{1}^{T}\mathbf{d}}{n}.$ (1)
Feld showed the following result, which may be proved by applying standard arguments in statistics [13, 38] or linear algebra [20].
Theorem 2.1 (Global friendship paradox [13]).
The inequality
$\frac{\mathbf{d}^{T}\mathbf{d}}{\mathbf{1}^{T}\mathbf{d}}-\frac{\mathbf{1}^{T}\mathbf{d}}{n}\geq 0$ (2)
always holds, with equality if and only if the network is regular.
Theorem 2.1 may interpreted as saying that, on average our friends have at least as many friends as us, with equality occuring only if we all have the same number of friends. Because this inequality deals with a node-wide average, and to distinguish it from (4) below, we will refer to (2) as the global friendship inequality.
Eom and Jo [11] considered the case where each node $i$ in the graph has an associated attribute, $x_{i}>0$. An appropriate generalisation of the inequality (2) is then
$\frac{\mathbf{d}^{T}\mathbf{x}}{\mathbf{1}^{T}\mathbf{d}}-\frac{\mathbf{1}^{T}\mathbf{x}}{n}\geq 0.$ (3)
We will refer to (3) as a generalized global friendship inequality; if this holds then, on average, our neighbours have more of the attribute $\mathbf{x}$ than we do. We emphasize that in the general case where the attribute vector $\mathbf{x}$ is independent of the network structure, there is no way of knowing a priori whether the inequality (3) will hold. However, there is a simple and intuitive characterisation: Eom and Jo [11] showed that the inequality is equivalent to the statement that $\mathbf{x}$ is nonnegatively correlated with the degree vector.
In [20] Higham considered the case where the attribute $\mathbf{x}$ is a network centrality measure, and hence is directly dependent on the network structure. Using classical eigenvector centrality [5, 6, 39], the importance of node $i$ can be measured by $x_{i}$, where $\mathbf{x}$ is the Perron-Frobenius eigenvector of $A$. Here, $\mathbf{x}$ has positive elements, $A\mathbf{x}$ is proportional to $\mathbf{x}$, and we assume without loss of generality the normalization $\mathbf{1}^{T}\mathbf{x}=1$. We then have the following result.
Theorem 2.2 (Global eigenvector centrality paradox [20]).
The generalized global friendship inequality (3) holds for eigenvector centrality, with equality if and only if the network is regular.
Theorem 2.2 may be interpreted as saying that, on average, our friends are at least as important as us, with equality occurring only if we all have the same number of friends.
2.2 Local Averages
The original article [13] mentioned an alternative, more localized, way to compare a node with its neighbours. For node $i$, we may record how the nodal degree $d_{i}$ compares with the average degree of its neighbours by computing
$\Delta_{i}=\frac{1}{d_{i}}\sum_{j=1}^{n}a_{ij}d_{j}-d_{i}.$
After making these local comparisons, we may then move to a graph-wise summary by asking whether the arithmetic average of the $\Delta_{i}$ values is nonnegative. Since $\mbox{$\Delta$}=D^{-1}A\mathbf{d}-\mathbf{d}$, we may write this inequality as
$\mathbf{1}^{T}(D^{-1}A\mathbf{d}-\mathbf{d})\geq 0,$ (4)
giving another sense in which, on average, our friends might have at least as many friends as us. We will refer to (4) as the local friendship inequality. In [13] Feld defined this concept and reported results for specific networks; however, unlike in the global case (2), he did not study it in general. Thirty years later, Cantwell et al. [7] proved the analogue of Theorem 2.1:
Theorem 2.3 (Local friendship paradox [7]).
The inequality (4) always holds, with equality if and only if the network is regular.
Cantwell et al. [7] also introduced the generalized version of (4) where degree is replaced by an attribute vector $\mathbf{x}$. Here, instead of the global version (3), we have
$\mathbf{1}^{T}(D^{-1}A\mathbf{x}-\mathbf{x})\geq 0,$ (5)
which we refer to as a generalized local friendship inequality. The authors in [7] showed that (5) is equivalent to the statement that $\mathbf{x}$ correlates nonnegatively with $\kappa$, where
$\kappa_{i}=\sum_{j=1}^{n}\frac{a_{ij}}{d_{j}}.$ (6)
Hazra and Verbitskiy [19] showed that the generalized local friendship inequality holds for eigenvector centrality and also for walk centrality, $\mathbf{x}=A^{\ell}\mathbf{1}$, where $\ell>1$ is an integer, and for Katz centrality [25], $(I-\alpha)\mathbf{x}=\mathbf{1}$, where $0<\alpha<1/\rho(A)$, with $\rho(A)$ denoting the spectral radius of $A$.
Theorem 2.4 (Generalized local friendship paradox for eigenvector, walk and Katz centralities [19]).
The generalized local friendship inequality (5) holds for eigenvector, walk and Katz centrality, with equality if and only if the network is regular.
3 Generalized Local Friendship Inequality for Nonbacktracking Eigenvector Centrality
For some networks, the standard eigenvector centrality measure has been observed to exhibit localization—attributing large values to a small subset of very central nodes, and failing to distinguish adequately between the remainder [1, 27, 30, 35]. An alternative nonbacktracking version of eigenvector centrality was proposed in [30].
The nonbacktracking eigenvector centrality vector $\mathbf{v}_{a}\in\mathbb{R}^{n}$ can be computed via the $2n$ by $2n$ eigenvalue problem
$\left[\begin{array}[]{cc}A&I-D\\ I&0\end{array}\right]\left[\begin{array}[]{c}\mathbf{v}_{a}\\ \mathbf{v}_{b}\end{array}\right]=\lambda\left[\begin{array}[]{c}\mathbf{v}_{a}\\ \mathbf{v}_{b}\end{array}\right].$ (7)
Here, it is known that a dominant eigenvalue $\lambda$ exists, that is real-valued and positive, and in the corresponding eigenvector both $\mathbf{v}_{a}$ and $\mathbf{v}_{b}$ can be chosen to have nonnegative entries [17, 30]. As suggested in [30], we may then take $\mathbf{v}_{a}$ to be the nonbacktracking eigenvector centrality measure. We will normalize $\mathbf{v}_{a}$ so that $\mathbf{1}^{T}\mathbf{v}_{a}=1$.
The eigenvalue problem (7) is intimately connected with the Hashimoto, or nonbacktracking, matrix [30], and also with nonbacktracking Katz-style centrality [17, Theorem 10.2].
We now show that the local version of the generalized Friendship inequality applies for this centrality measure.
Theorem 3.1 (Local nonbacktracking eigenvector centrality paradox).
The generalized local friendship inequality (5) holds for nonbacktracking eigenvector centrality.
Proof.
By construction, we have
$A\mathbf{v}_{a}=\lambda\mathbf{v}_{a}+(D-I)\mathbf{v}_{b}.$ (8)
Premultiplying by $\mathbf{1}^{T}$, and rearranging, gives
$\lambda=\mathbf{d}^{T}\mathbf{v}_{a}-\mathbf{1}^{T}(D-I)\mathbf{v}_{b}.$ (9)
Starting from the left-hand side of the inequality (5) with $\mathbf{x}=\mathbf{v}_{a}$, we have, using (8),
$\mathbf{1}^{T}\left(D^{-1}A\mathbf{v}_{a}-\mathbf{v}_{a}\right)$ $=$ $\mathbf{1}^{T}\left(D^{-1}(\lambda\mathbf{v}_{a}+(D-I)\mathbf{v}_{b})-\mathbf{v}_{a}\right)$
$=$ $\lambda\mathbf{d}^{-T}\mathbf{v}_{a}+\mathbf{1}^{T}(D-I)\mathbf{v}_{b}-1.$
Using the expression (9) for $\lambda$, this gives
$\mathbf{1}^{T}\left(D^{-1}A\mathbf{v}_{a}-\mathbf{v}_{a}\right)$ $=$ $\left(\mathbf{d}^{T}\mathbf{v}_{a}-\mathbf{1}^{T}(D-I)\mathbf{v}_{b}\right)\mathbf{d}^{-T}\mathbf{v}_{a}+\mathbf{1}^{T}(D-I)\mathbf{v}_{b}-1$ (10)
$=$ $(\mathbf{d}^{T}\mathbf{v}_{a})(\mathbf{d}^{-T}\mathbf{v}_{a})-1+\mathbf{1}^{T}(D-I)\mathbf{v}_{b}\left(1-\mathbf{d}^{-T}\mathbf{v}_{a}\right).$
Now the arithmetic mean-harmonic mean inequality shows that $(\mathbf{d}^{T}\mathbf{v}_{a})(\mathbf{d}^{-T}\mathbf{v}_{a})\geq 1$. Because $\mathbf{1}$, DID-I and $\mathbf{v}_{b}$ contain nonnegative elements, we have $\mathbf{1}^{T}(D-I)\mathbf{v}_{b}\geq 0$. Finally, since $\mathbf{1}^{T}\mathbf{v}_{a}=1$ and each $d_{i}\geq 1$, we have $\mathbf{d}^{-T}\mathbf{v}_{a}\leq 1$. So the right-hand side of (10) is nonnegative, as required. ∎
We note that in the statement of Theorem 3.1 we cannot add that equality holds if and only if the network is regular. To see this, suppose $\mathbf{d}=A\mathbf{1}=d\mathbf{1}$ for some common degree $d\geq 2$. Then the the right hand side of (10) reduces to $(d-1)(1-\frac{1}{d})\mathbf{1}^{T}\mathbf{v}_{b}$, which is nonzero.
4 Generalized Friendship Paradoxes: Global versus Local
The generalized global and local friendship inequalities are similar in the sense that both hold universally for the attributes of degree (Theorems 2.1 and 2.3) and eigenvector centrality (Theorems 2.2 and 2.4). However, the characterizations mentioned in subsections 2.1 and 2.2 suggest that they are not equivalent—the global version corresponds to correlation with the degree, while the local version corresponds to correlation with the vector $\kappa$ in (6). Here, we confirm with a concrete example that all four possibilities of the two inequalities holding/not holding may arise. We consider the four-node network with adjacency matrix
$A=\left[\begin{array}[]{cccc}0&1&0&0\\ 1&0&1&1\\ 0&1&0&1\\ 0&1&1&0\end{array}\right].$
Given an attribute vector $\mathbf{x}$ it is straightforward to show that the generalized global friendship inequality (3) and generalized local friendship inequality (5) are equivalent to
$x_{2}-x_{1}>0\quad\text{and}\quad-4x_{1}+6x_{2}-x_{3}-x_{4}>0,$
respectively. It follows that
  • with $x_{1}=1$, $x_{2}=2$ and $x_{3}=x_{4}=6$ we have (3) but not (5),
  • with $x_{1}=7$, $x_{2}=6$ and $x_{3}=x_{4}=2$ we have (5) but not (3),
  • with $x_{1}=2$, $x_{2}=1$ and $x_{3}=x_{4}=1$ we have neither (3) nor (5),
and (immediately from Theorems 2.1 and 2.3) with $\mathbf{x}=\mathbf{d}$ we have both (3) and (5).
5 A Reverse Friendship Paradox
By analogy with (3) we may say that a reverse generalized global friendship inequality arises if, on average, a node has more of an attribute than its neighbours; that is,
$\frac{\mathbf{d}^{T}\mathbf{x}}{\mathbf{1}^{T}\mathbf{d}}-\frac{\mathbf{1}^{T}\mathbf{x}}{n}\leq 0.$ (11)
Here we show the intuitively reasonable result that the reciprocal of degree has this property. Thus if we regard $\mathbf{d}^{-1}$ as a measure of loneliness, then on average we are at least as lonely as our friends.
Theorem 5.1 (Global loneliness paradox).
The reverse generalized global friendship inequality (11) holds for $\mathbf{x}=\mathbf{d}^{-1}$, with equality if and only if the network is regular.
Proof.
In this case (11) becomes
$\frac{\mathbf{d}^{T}\mathbf{d}^{-1}}{\mathbf{1}^{T}\mathbf{d}}-\frac{\mathbf{1}^{T}\mathbf{d}^{-1}}{n}\leq 0,$
which rearranges to
$\frac{n}{\mathbf{1}^{T}\mathbf{d}}-\frac{\mathbf{1}^{T}\mathbf{d}^{-1}}{n}\leq 0.$
The result now follows directly from the arithmetic mean-harmonic mean inequality [31]. ∎
6 A Geometric Mean Friendship Paradox
In this section we will make use of the following lemma.
Lemma 6.1.
Given a strictly monotonically increasing function $f:\mathbb{R}\to\mathbb{R}$, the generalized global friendship inequality (3) holds for $x_{i}=f(d_{i})$, with equality if and only if the network is regular.
Proof.
Since $f$ is monotonically increasing, we have, for any $i$ and $j$,
$\left(d_{i}-d_{j}\right)\left(f(d_{i})-f(d_{j})\right)\geq 0,$
with equality if and only if $d_{i}=d_{j}$. Summing, we obtain,
$\sum_{i=1}^{n}\sum_{j=1}^{n}\left(d_{i}-d_{j}\right)\left(f(d_{i})-f(d_{j})\right)\geq 0,$
with equality if and only if the network is regular. Upon simplification and rearrangement, this inequality reduces to
$\sum_{i=1}^{n}\frac{d_{i}f(d_{i})}{\mathbf{1}^{T}\mathbf{d}}-\frac{1}{n}\sum_{i=1}^{n}f(d_{i})\geq 0,$
as required. ∎
As we mentioned in section 2, Eom and Jo [11] showed that (3) is equivalent to asking for $\mathbf{x}$ to be nonnegatively correlated with the degree vector. Lemma 6.1 shows the intuitively reasonable result that applying a monotonic function to the degree vector respects the inequality.
The aim of this section is to consider using the geometric mean in (2) rather than the arithmetic mean. As mentioned in the start of subsection 2.1, the friend-of-friend list includes each degree value $d_{i}$ exactly $d_{i}$ times, and the list involves $\mathbf{1}^{T}\mathbf{d}$ numbers. So the friend-of-friend geometric mean is
$\left(\prod_{i=1}^{n}{d_{i}}^{d_{i}}\right)^{1/\mathbf{1}^{T}\mathbf{d}}.$
The friend geometric mean is
$\left(\prod_{i=1}^{n}{d_{i}}\right)^{1/n}.$
So we have a global geometric mean friendship paradox if
$\left(\prod_{i=1}^{n}{d_{i}}^{d_{i}}\right)^{1/\mathbf{1}^{T}\mathbf{d}}\geq\left(\prod_{i=1}^{n}{d_{i}}\right)^{1/n}.$ (12)
Theorem 6.1 (Global geometric mean friendship paradox).
The global geometric mean friendship inequality (12) holds, with equality if and only if the network is regular.
Proof.
Taking logarithms in (12) gives
$\frac{1}{\mathbf{1}^{T}\mathbf{d}}\sum_{i=1}^{n}\log{d_{i}}^{d_{i}}\geq\frac{1}{n}\sum_{i=1}^{n}\log{d_{i}},$
which is equivalent to
$\sum_{i=1}^{n}\frac{d_{i}\log{d_{i}}}{\mathbf{1}^{T}\mathbf{d}}\geq\frac{1}{n}\sum_{i=1}^{n}\log{d_{i}}.$ (13)
The result now follows from Lemma 6.1.
Acknowledgement
DJH was supported by a Leverhulme Fellowship. FT was partially funded by the PRIN-MUR project MOLE code 2022ZK5ME7 and by PRIN-PNRR project FIN4GEO CUP P2022BNB97.