### v2

# Generalized Friendship Paradoxes in Network Science

Abstract
Generalized friendship paradoxes occur when, on average, our friends have more of some attribute than us. These paradoxes are relevant to many aspects of human interaction, notably in social science and epidemiology. Here, we derive new theoretical results concerning the inevitability of a paradox arising, using a linear algebra perspective. Following the seminal 1991 work of Scott L. Feld, we consider two distinct ways to measure and compare averages, which may be regarded as global and local. For global averaging, we show that a generalized friendship paradox holds for a large family of walk-based centralities, including Katz centrality and total subgraph communicability, and also for nonbacktracking eigenvector centrality. Defining loneliness as the reciprocal of the number of friends, we show that for this attribute the generalized friendship paradox always holds in reverse. In this sense, we are always more lonely, on average, than our friends. We also derive a global averaging paradox result for the case where the arithmetic mean is replaced by the geometric mean. The nonbacktracking eigenvector centrality paradox is also established for the case of local averaging.
1 Motivation
In 1991, the sociologist Scott L. Feld pointed out a sampling bias phenomenon that applies to any unweighted, undirected network [17]. In the context of social networks, on average, our friends have more friends than we do. This effect has subsequently become known as the Friendship Paradox. With regard to our online interactions, it has been argued that the effect distorts our perceptions [30] and contributes to our dissatisfaction [6, 13]. On a more positive note, the paradox may be exploited in the design of strategies to monitor or control the spread of information or disease [5, 11, 10, 19, 23, 36, 37, 45], including vaccination programmes. More generally, it may be used to sample more efficiently from the tail of a degree distribution [42].
In [14] Eom and Jo looked at a Generalized Friendship Paradox: given an externally derived attribute, on average, do our neighbours have more of this attribute than we do? For example, in a scientific collaboration network, do our coauthors have a higher citation count than us on average? Such issues have subsequently been studied in a range of empirical tests, leading to concepts such as the activity and virality paradoxes [28], the H-index paradox [2], the happiness paradox [45] and the enmity paradox [21]. A further direction of research focuses attention on specific classes of networks, as in [9, 16, 31, 32, 47].
The original work of Feld [17] set up two distinct types of averaging process. Following [21] we will distinguish between them by using the terms global and local. (We note that the terms “ego” and “alter” are used in [36].) Among the empirical tests for paradox effects across various application domains, some emphasize local averaging [2, 12, 6, 9, 34, 40, 53, 54] and others global averaging [25, 28, 41]. In this work we contribute to the understanding of both types of generalized friendship paradox by highlighting new cases where a paradox is guaranteed to hold for any network. We thereby add to the theoretical inevitability results on global averaging in [17, 27] and on local averaging in [9, 14, 26]. We make several contributions:
  • Theorem 3.1 and Corollary 3.1 show that a global averaging paradox holds for a wide class of walk-based centrality measures.
  • Theorems 3.2 and 4.1 show that both global and local averaging paradoxes hold for a nonbacktracking version of eigenvector centrality.
  • Theorems 6.1 and 6.2 show that both global and local averaging paradoxes hold for loneliness (defined as the reciprocal of degree).
  • Theorem 7.1 shows that a global geometric mean paradox holds for degree.
In deriving results, to minimize confusion we will use the phrase “inequality” for expressions that arise when we compare related quantities and reserve the phrase “paradox” for a result showing that the inequality holds for all networks.
2 Definitions and Existing Results
We consider undirected, unweighted, connected networks with $n$ nodes and no self-loops. We let $A\in\mathbb{R}^{n\times n}$ denote the adjacency matrix. So $A$ is a symmetric binary matrix with zero diagonal such that $a_{ij}=1$ if nodes $i$ and $j$ are connected and $a_{ij}=0$ otherwise.
We define the degree vector $\mathbf{d}\in\mathbb{R}^{n}$ by
$\mathbf{d}=A\mathbf{1},$
where $\mathbf{1}\in\mathbb{R}^{n}$ is the vector of ones. So $d_{i}$ is the degree of node $i$. In a social network context, $d_{i}$ measures the number of friends that node $i$ possesses. We use $\mathbf{d}^{-1}$ to denote the vector with $i$th element $1/d_{i}$; so $\mathbf{d}^{-1}$ contains the reciprocal degrees. We use $\mathbf{d}^{-T}$ to denote the transpose of $\mathbf{d}^{-1}$. Hence, when $\mathbf{x}\in\mathbb{R}^{n}$ we have $\mathbf{d}^{-T}\mathbf{x}=\sum_{i=1}^{n}x_{i}/d_{i}$. We let $D\in\mathbb{R}^{n\times n}$ denote the diagonal degree matrix, for which $D_{ii}=d_{i}$.
2.1 Global Averages
By definition, the (arithmetic) average degree, or average number of friends, taken over all nodes is given by $\mathbf{1}^{T}\mathbf{d}/n$. Suppose instead that we consider all nodes and compute the average number of friends-of-friends. In this computation, every node appears as a friend $d_{i}$ times, and each time contributes its $d_{i}$ friends to the sum, giving a total of $\mathbf{d}^{T}\mathbf{d}$. The number of terms in this sum is $\mathbf{1}^{T}\mathbf{d}$, since every edge contributes twice. It follows that the difference between the friend-of-friend average and the friend average is
$\frac{\mathbf{d}^{T}\mathbf{d}}{\mathbf{1}^{T}\mathbf{d}}-\frac{\mathbf{1}^{T}\mathbf{d}}{n}.$ (1)
Feld showed the following result, which may be proved by applying standard arguments in statistics [17, 49] or linear algebra [27].
Theorem 2.1 (Global friendship paradox [17]).
The inequality
$\frac{\mathbf{d}^{T}\mathbf{d}}{\mathbf{1}^{T}\mathbf{d}}-\frac{\mathbf{1}^{T}\mathbf{d}}{n}\geq 0$ (2)
always holds, with equality if and only if the network is regular.
Theorem 2.1 may interpreted as saying that, on average our friends have at least as many friends as us, with equality occuring only if we all have the same number of friends. Because this inequality deals with a node-wide average, and to distinguish it from (4) below, we will refer to (2) as the global friendship inequality.
Eom and Jo [14] considered the case where each node $i$ in the graph has an associated attribute, $x_{i}>0$. An appropriate generalisation of the inequality (2) is then
$\frac{\mathbf{d}^{T}\mathbf{x}}{\mathbf{1}^{T}\mathbf{d}}-\frac{\mathbf{1}^{T}\mathbf{x}}{n}\geq 0.$ (3)
We will refer to (3) as a generalized global friendship inequality; if this holds then, on average, our neighbours have more of the attribute $\mathbf{x}$ than we do. We emphasize that in the general case where the attribute vector $\mathbf{x}$ is independent of the network structure, there is no way of knowing a priori whether the inequality (3) will hold. However, there is a simple and intuitive characterisation: Eom and Jo [14] showed that the inequality is equivalent to the statement that $\mathbf{x}$ is nonnegatively correlated with the degree vector.
The work of Higham [27] was the first to consider the case where the attribute $\mathbf{x}$ is a network centrality measure (other than degree), and hence is directly dependent on the network structure. According to classical eigenvector centrality [7, 8, 52], the importance of node $i$ can be measured by $x_{i}$, where $\mathbf{x}$ is the Perron-Frobenius eigenvector of $A$. Here, $\mathbf{x}$ has positive elements, $A\mathbf{x}$ is proportional to $\mathbf{x}$, and we assume without loss of generality the normalization $\mathbf{1}^{T}\mathbf{x}=1$. We then have the following result.
Theorem 2.2 (Global eigenvector centrality paradox [27]).
The generalized global friendship inequality (3) holds for eigenvector centrality, with equality if and only if the network is regular.
Theorem 2.2 may be interpreted as saying that, on average, our friends are at least as important as us, with equality occurring only if we all have the same number of friends. Results of this type add further weight to the arguments proposed in the social sciences that friendship paradox effects contribute to the perceived levels of dissatisfaction among those who engage in online interaction. They also support the empirical findings that the paradox can be exploited for monitoring or controlling the spread of information or disease.
2.2 Local Averages
The original article [17] mentioned an alternative, more localized, way to compare a node with its neighbours. For node $i$, we may record how the nodal degree $d_{i}$ compares with the average degree of its neighbours by computing
$\Delta_{i}=\frac{1}{d_{i}}\sum_{j=1}^{n}a_{ij}d_{j}-d_{i}.$
After making these local comparisons, we may then move to a graph-wise summary by asking whether the arithmetic average of the $\Delta_{i}$ values is nonnegative. Since $\mbox{$\Delta$}=D^{-1}A\mathbf{d}-\mathbf{d}$, we may write this inequality as
$\mathbf{1}^{T}(D^{-1}A\mathbf{d}-\mathbf{d})\geq 0,$ (4)
giving another sense in which, on average, our friends might have at least as many friends as us. We will refer to (4) as the local friendship inequality. In [17] Feld defined this concept and reported results for specific networks; however, unlike in the global case (2), he did not study it in general. Thirty years later, Cantwell et al. [9] proved the analogue of Theorem 2.1:
Theorem 2.3 (Local friendship paradox [9]).
The inequality (4) always holds, with equality if and only if the network is regular.
Cantwell et al. [9] also introduced the generalized version of (4) where degree is replaced by an attribute vector $\mathbf{x}$. Here, instead of the global version (3), we have
$\mathbf{1}^{T}(D^{-1}A\mathbf{x}-\mathbf{x})\geq 0,$ (5)
which we refer to as a generalized local friendship inequality. The authors in [9] showed that (5) is equivalent to the statement that $\mathbf{x}$ correlates nonnegatively with $\kappa$, where
$\kappa_{i}=\sum_{j=1}^{n}\frac{a_{ij}}{d_{j}}.$ (6)
As indicated in the theorem below, Hazra and Verbitskiy [26] showed that the generalized local friendship inequality holds for eigenvector centrality and also for walk centrality, defined as
$\mathbf{x}=A^{\ell}\mathbf{1},$ (7)
where $\ell>1$ is an integer, and for Katz centrality [33], defined as
$(I-\alpha)\mathbf{x}=\mathbf{1},$ (8)
where $0<\alpha<1/\rho(A)$, with $\rho(A)$ denoting the spectral radius of $A$.
Theorem 2.4 (Generalized local friendship paradox for eigenvector, walk and Katz centralities [26]).
The generalized local friendship inequality (5) holds for eigenvector, walk and Katz centrality, with equality if and only if the network is regular.
We also note that the analysis in [26] establishes equality if and only if the network is regular in the cases of eigenvector and Katz centrality.
In the following two sections we demonstrate that further global and local paradoxes may be established.
3 Generalized Global Friendship Paradoxes
In this section we focus on the generalized global friendship inequality (3) and derive new results concerning centrality measures.
3.1 Generalized Global Friendship Paradox for Katz-style Centralities
We begin with a wide class of matrix-function based centralities; see, for example, [15] for further motivation.
Theorem 3.1 (Generalized global friendship paradox for $f(A)\mathbf{1}$).
The generalized global friendship inequality (3) holds for $\mathbf{x}=f(A)\mathbf{1}$, where $f$ is an increasing function defined over the spectrum of $A$. (In particular, it holds when $f$ has a convergent power series with positive coefficients.) Furthermore, if $f$ is strictly increasing then equality holds if and only if the associated graph is regular.
Proof.
Since $A$ is real and symmetric, it admits an orthonormal eigenbasis $\{\mathbf{v}_{j}\}$ with corresponding real eigenvalues $\lambda_{j}$. So we may write (with all sums running from $1$ to $n$)
$A=\sum_{j}\lambda_{j}\mathbf{v}_{j}\mathbf{v}_{j}^{T},\qquad\mathbf{1}=\sum_{j}(\mathbf{v}_{j}^{T}\mathbf{1})\,\mathbf{v}_{j},$ (9)
and also
$\mathbf{x}=f(A)\mathbf{1}=\sum_{j}f(\lambda_{j})(\mathbf{v}_{j}^{T}\mathbf{1})\mathbf{v}_{j},\qquad\mathbf{d}=A\mathbf{1}=\sum_{j}\lambda_{j}(\mathbf{v}_{j}^{T}\mathbf{1})\,\mathbf{v}_{j}.$ (10)
Letting $w_{j}:=(\mathbf{v}_{j}^{T}\mathbf{1})^{2}\geq 0$ it follows that
$\mathbf{1}^{T}\mathbf{d}=\sum_{j}\lambda_{j}w_{j},\qquad\mathbf{d}^{T}\mathbf{x}=\sum_{j}\lambda_{j}f(\lambda_{j})w_{j},\qquad\mathbf{1}^{T}\mathbf{x}=\sum_{j}f(\lambda_{j})w_{j}.$
The inequality (3) is therefore equivalent to
$n\sum_{j}\lambda_{j}f(\lambda_{j})w_{j}\;\geq\;\Big(\sum_{j}\lambda_{j}w_{j}\Big)\Big(\sum_{j}f(\lambda_{j})w_{j}\Big).$ (11)
Now define the quantity
$\mu:=\frac{1}{n}\sum_{j}\lambda_{j}w_{j}=\frac{\mathbf{1}^{T}\mathbf{d}}{n},$
so that the desired inequality (11) may be written
$\sum_{j}w_{j}f(\lambda_{j})(\lambda_{j}-\mu)\geq 0.$ (12)
To make progress, we will establish the identity
$\sum_{j}w_{j}f(\lambda_{j})(\lambda_{j}-\mu)=\frac{1}{2n}\sum_{i,j}w_{i}w_{j}(\lambda_{i}-\lambda_{j})(f(\lambda_{i})-f(\lambda_{j})).$ (13)
Using the fact that
$\sum_{i}w_{i}=\sum_{i}(\mathbf{v}_{i}^{T}\mathbf{1})^{2}=\|\mathbf{1}\|_{2}^{2}=n,$
we have
$\frac{1}{2n}\sum_{i,j}w_{i}w_{j}(\lambda_{i}-\lambda_{j})(f(\lambda_{i})-f(\lambda_{j}))$ $=$ $\frac{1}{2n}\sum_{i,j}w_{i}w_{j}\Big(\lambda_{i}f(\lambda_{i})-\lambda_{i}f(\lambda_{j})-\lambda_{j}f(\lambda_{i})+\lambda_{j}f(\lambda_{j})\big)$
$=$ $\frac{1}{2n}\Big(\Big(\sum_{j}w_{j}\Big)\Big(\sum_{i}w_{i}\lambda_{i}f(\lambda_{i})\Big)+\Big(\sum_{i}w_{i}\Big)\Big(\sum_{j}w_{j}\lambda_{j}f(\lambda_{j})\Big)$
$\mbox{}-\sum_{i,j}w_{i}w_{j}\lambda_{i}f(\lambda_{j})-\sum_{i,j}w_{i}w_{j}\lambda_{j}f(\lambda_{i})\Big)$
$=$ $\frac{1}{2n}\Big(2n\sum_{i}w_{i}\lambda_{i}f(\lambda_{i})-2\sum_{i,j}w_{i}w_{j}\lambda_{i}f(\lambda_{j})\Big),$
$=$ $\sum_{i}w_{i}\lambda_{i}f(\lambda_{i})-\frac{1}{n}\Big(\sum_{i}w_{i}\lambda_{i}\Big)\Big(\sum_{j}w_{j}f(\lambda_{j})\Big),$
$=$ $\sum_{i}w_{i}\lambda_{i}f(\lambda_{i})-\mu\sum_{j}w_{j}f(\lambda_{j}),$
$=$ $\sum_{i}w_{i}f(\lambda_{i})(\lambda_{j}-\mu),$
as required.
Because $f$ is increasing we have $(\lambda_{i}-\lambda_{j})(f(\lambda_{i})-f(\lambda_{j}))q\geq 0$ for any $i$ and $j$. Summing, we see that the expression on the right hand side of (13) is greater than or equal to zero, and hence (12) holds.
If $f$ is strictly increasing then it is clear from the identity (13) that equality holds and only if all $\lambda_{i}$ are equal for eigenvalues whose eigenvector is orthogonal to $\mathbf{1}$. In this case, from (9) and (10), we have
$\mathbf{d}=\sum_{j}\lambda_{j}(\mathbf{v}_{j}^{T}\mathbf{1})\mathbf{v}_{j}=\lambda\sum_{j}(\mathbf{v}_{j}^{T}\mathbf{1})\mathbf{v}_{i}=\lambda\mathbf{1},$
so the graph is regular. On the other hand, if the graph is regular, then
$\frac{\mathbf{d}^{T}\mathbf{x}}{\mathbf{1}^{T}\mathbf{d}}=\frac{\lambda\mathbf{1}^{T}\mathbf{x}}{\lambda\mathbf{1}^{T}\mathbf{1}}=\frac{\|\mathbf{x}\|_{1}}{n},$
completing the proof. ∎
Theorem 3.1 immediately covers walk centrality (7), Katz centrality (8) and total subgraph communicability [3] (also known as exponential centrality), which is defined by
$\mathbf{x}=\exp(A)\mathbf{1}.$ (14)
Corollary 3.1 (Examples arising from $f(A)\mathbf{1}$).
The generalized global friendship inequality (3) holds for walk centrality (7) when $\ell$ is odd, Katz centrality (8) and total subgraph communicability (14). Furthermore, in these cases equality holds if and only if the associated graph is regular.
3.2 Generalized Global Friendship Paradox for Nonbacktracking Eigenvector Centrality
For some networks, the standard eigenvector centrality measure has been observed to exhibit localization—attributing large values to a small subset of very central nodes, and failing to distinguish adequately between the remainder [1, 35, 38, 44]. An alternative nonbacktracking version of eigenvector centrality was proposed in [38].
The nonbacktracking eigenvector centrality vector $\mathbf{v}_{a}\in\mathbb{R}^{n}$ can be computed via the $2n$ by $2n$ eigenvalue problem
$\left[\begin{array}[]{cc}A&I-D\\ I&0\end{array}\right]\left[\begin{array}[]{c}\mathbf{v}_{a}\\ \mathbf{v}_{b}\end{array}\right]=\lambda\left[\begin{array}[]{c}\mathbf{v}_{a}\\ \mathbf{v}_{b}\end{array}\right].$ (15)
Here, it is known that a dominant eigenvalue $\lambda$ of this $2n$ by $2n$ matrix exists, that is real-valued and positive, and in the corresponding eigenvector both $\mathbf{v}_{a}$ and $\mathbf{v}_{b}$ can be chosen to have nonnegative entries [24, 38]. As suggested in [38], we may then take $\mathbf{v}_{a}$ to be the nonbacktracking eigenvector centrality measure. We will normalize $\mathbf{v}_{a}$ so that $\mathbf{1}^{T}\mathbf{v}_{a}=1$.
The eigenvalue problem (15) is intimately connected with the nonbacktracking, or Hashimoto, matrix [38, 43, 48] and also with nonbacktracking Katz-style centrality [24, Theorem 10.2].
We now show that a generalized global friendship paradox holds for the nonbacktracking eigenvalue centrality measure.
Theorem 3.2 (Generalized global nonbacktracking eigenvector centrality paradox).
The generalized global friendship inequality (3) holds for nonbacktracking eigenvector centrality (defined via $\mathbf{x}=\mathbf{v}_{a}$ in (15)).
Proof.
First we note that $\lambda\geq 1$, since
$\left[\begin{array}[]{cc}A&I-D\\ I&0\end{array}\right]\left[\begin{array}[]{c}\mathbf{1}\\ \mathbf{1}\end{array}\right]=\left[\begin{array}[]{c}\mathbf{1}\\ \mathbf{1}\end{array}\right].$
If $\lambda=1$ then $\mathbf{v}_{a}=\mathbf{1}$ and the inequality holds with equality.
Hence, we now assume that $\lambda>1$.
From (15), we have
$A\mathbf{v}_{a}=\lambda\mathbf{v}_{a}+(D-I)\mathbf{v}_{b},$ (16)
and
$\mathbf{v}_{a}=\lambda\mathbf{v}_{b}.$
So
$A\mathbf{v}_{a}=\lambda\mathbf{v}_{a}+(D-I)\frac{\mathbf{v}_{a}}{\lambda},$
which may be written
$\lambda A\mathbf{v}_{a}=((\lambda^{2}-1)I+D)\mathbf{v}_{a}.$
So
$\lambda\mathbf{1}^{T}A\mathbf{v}_{a}=\mathbf{1}^{T}((\lambda^{2}-1)I+D)\mathbf{v}_{a}.$
Using $\mathbf{1}^{T}\mathbf{v}_{a}=1$, this simplifies to
$(\lambda-1)\mathbf{d}^{T}\mathbf{v}_{a}=(\lambda^{2}-1).$
Equivalently
$\mathbf{d}^{T}\mathbf{v}_{a}=\lambda+1.$
So (3) holds for $\mathbf{x}=\mathbf{v}_{a}$ if and only if
$\lambda\geq\frac{\mathbf{1}^{T}\mathbf{d}}{n}-1.$ (17)
We now make use of the connection between the eigenvalue problem (15), the nonbacktracking matrix and the radius of convergence, $R$, of the Ihara zeta function of the graph. It is known [29, Thm 5.6.15 applied to the matrix $zA$] that $\lambda$ is equal to the largest eigenvalue of the nonbacktracking matrix, which itself is equal to the reciprocal of $R$; so $\lambda=1/R$. Terras [50] conjectured that
$\frac{1}{R}\geq\frac{\mathbf{1}^{T}\mathbf{d}}{n}-1$
(which is precisely (17)) for a graph that does not have a node of degree one and is not a cycle. This conjecture was proved by Saito in 2018 [46].
To finish we need to deal with the remaining cases where the graph has a node of degree one or the graph is a cycle.
It is shown in [22, Corollary 3.2.1] that increasing the size of the graph by taking one node and joining it to a tree does not change $\lambda$. Hence, removing a single node of degree one, along with the corresponding edge, does not change $\lambda$. See also [43, Theorem 1.2] for a more direct argument that also applies to directed graphs. We may therefore iteratively remove all nodes of degree one until we have a graph where all nodes have degree at least two, whence the analysis above applies if the remaining graph is not a cycle.
If the graph is a cycle, then from [51, Proposition 3.6] we have $\lambda=1$, and hence, as argued at the start of the proof, equality arises. ∎
4 Generalized Local Friendship Paradoxes
We now prove the equivalent of Theorem 3.2 for the local case.
Theorem 4.1 (Generalized local nonbacktracking eigenvector centrality paradox).
The generalized local friendship inequality (5) holds for nonbacktracking eigenvector centrality (defined via $\mathbf{x}=\mathbf{v}_{a}$ in (15)).
Proof.
Premultiplying (16) by $\mathbf{1}^{T}$, using $\mathbf{1}^{T}\mathbf{v}_{a}=1$ and rearranging, we have
$\lambda=\mathbf{d}^{T}\mathbf{v}_{a}-\mathbf{1}^{T}(D-I)\mathbf{v}_{b}.$ (18)
Starting from the left-hand side of the inequality (5) with $\mathbf{x}=\mathbf{v}_{a}$, we have, using (16),
$\mathbf{1}^{T}\left(D^{-1}A\mathbf{v}_{a}-\mathbf{v}_{a}\right)$ $=$ $\mathbf{1}^{T}\left(D^{-1}(\lambda\mathbf{v}_{a}+(D-I)\mathbf{v}_{b})-\mathbf{v}_{a}\right)$
$=$ $\lambda\mathbf{d}^{-T}\mathbf{v}_{a}+\mathbf{1}^{T}(D-I)\mathbf{v}_{b}-1.$
Using the expression (18) for $\lambda$, this gives
$\mathbf{1}^{T}\left(D^{-1}A\mathbf{v}_{a}-\mathbf{v}_{a}\right)$ $=$ $\left(\mathbf{d}^{T}\mathbf{v}_{a}-\mathbf{1}^{T}(D-I)\mathbf{v}_{b}\right)\mathbf{d}^{-T}\mathbf{v}_{a}+\mathbf{1}^{T}(D-I)\mathbf{v}_{b}-1$ (19)
$=$ $(\mathbf{d}^{T}\mathbf{v}_{a})(\mathbf{d}^{-T}\mathbf{v}_{a})-1+\mathbf{1}^{T}(D-I)\mathbf{v}_{b}\left(1-\mathbf{d}^{-T}\mathbf{v}_{a}\right).$
Now the arithmetic mean-harmonic mean inequality shows that $(\mathbf{d}^{T}\mathbf{v}_{a})(\mathbf{d}^{-T}\mathbf{v}_{a})\geq 1$. Because $\mathbf{1}$, DID-I and $\mathbf{v}_{b}$ contain nonnegative elements, we have $\mathbf{1}^{T}(D-I)\mathbf{v}_{b}\geq 0$. Finally, since $\mathbf{1}^{T}\mathbf{v}_{a}=1$ and each $d_{i}\geq 1$, we have $\mathbf{d}^{-T}\mathbf{v}_{a}\leq 1$. So the right-hand side of (19) is nonnegative, as required. ∎
We note that in the statement of Theorem 4.1 we cannot add that equality holds if and only if the network is regular. To see this, suppose $\mathbf{d}=A\mathbf{1}=d\mathbf{1}$ for some common degree $d\geq 2$. Then the the right hand side of (19) reduces to $(d-1)(1-\frac{1}{d})\mathbf{1}^{T}\mathbf{v}_{b}$, which is nonzero.
5 Generalized Friendship Paradoxes: Global versus Local
The generalized global and local friendship inequalities are similar in the sense that both hold universally for the attributes of degree (Theorems 2.1 and 2.3) and eigenvector centrality (Theorems 2.2 and 2.4). However, the characterizations mentioned in subsections 2.1 and 2.2 suggest that they are not equivalent—the global version corresponds to correlation with the degree, while the local version corresponds to correlation with the vector $\kappa$ in (6). Here, we confirm with a concrete example that all four possibilities of the two inequalities holding/not holding may arise. We consider the four-node network with adjacency matrix
$A=\left[\begin{array}[]{cccc}0&1&0&0\\ 1&0&1&1\\ 0&1&0&1\\ 0&1&1&0\end{array}\right].$
Given an attribute vector $\mathbf{x}$ it is straightforward to show that the generalized global friendship inequality (3) and generalized local friendship inequality (5) are equivalent to
$x_{2}-x_{1}\geq 0\quad\text{and}\quad-4x_{1}+6x_{2}-x_{3}-x_{4}\geq 0,$
respectively. It follows that
  • with $x_{1}=1$, $x_{2}=2$ and $x_{3}=x_{4}=6$ we have (3) but not (5),
  • with $x_{1}=7$, $x_{2}=6$ and $x_{3}=x_{4}=2$ we have (5) but not (3),
  • with $x_{1}=2$, $x_{2}=1$ and $x_{3}=x_{4}=1$ we have neither (3) nor (5),
and (immediately from Theorems 2.1 and 2.3) with $\mathbf{x}=\mathbf{d}$ we have both (3) and (5).
6 Loneliness Paradoxes
By analogy with (3) we may say that a reverse generalized global friendship inequality arises if, on average, a node has more of an attribute than its neighbours; that is,
$\frac{\mathbf{d}^{T}\mathbf{x}}{\mathbf{1}^{T}\mathbf{d}}-\frac{\mathbf{1}^{T}\mathbf{x}}{n}\leq 0.$ (20)
Here we show the intuitively reasonable result that the reciprocal of degree has this property. Thus if we regard $\mathbf{d}^{-1}$ as a measure of loneliness, then on average we are at least as lonely as our friends.
Theorem 6.1 (Global loneliness paradox).
The reverse generalized global friendship inequality (20) holds for $\mathbf{x}=\mathbf{d}^{-1}$, with equality if and only if the network is regular.
Proof.
In this case (20) becomes
$\frac{\mathbf{d}^{T}\mathbf{d}^{-1}}{\mathbf{1}^{T}\mathbf{d}}-\frac{\mathbf{1}^{T}\mathbf{d}^{-1}}{n}\leq 0,$
which simplifies to
$\frac{n}{\mathbf{1}^{T}\mathbf{d}}-\frac{\mathbf{1}^{T}\mathbf{d}^{-1}}{n}\leq 0.$
The result now follows directly from the arithmetic mean-harmonic mean inequality [39]. ∎
Similarly, we may say that a reverse generalized local friendship inequality arises if (5) is reversed, so that
$\mathbf{1}^{T}(D^{-1}A\mathbf{x}-\mathbf{x})\leq 0.$ (21)
The next result shows that the corresponding paradox holds.
Theorem 6.2 (Local loneliness paradox).
The reverse generalized local friendship inequality (21) holds for $\mathbf{x}=\mathbf{d}^{-1}$, with equality if and only if the network is regular.
Proof.
We use similar arguments to those in [9, Section 2]. Taking $x_{i}=1/d_{i}$, in the left hand side of (21) we have
$\mathbf{1}^{T}(D^{-1}A\mathbf{x}-\mathbf{x})=\sum_{i,j}\left(\frac{a_{i,j}}{d_{i}d_{j}}\right)-\left(\sum_{i}\frac{1}{d_{i}}\right)=\sum_{i,j}\left(\frac{a_{i,j}}{d_{i}d_{j}}-\frac{a_{i,j}}{d_{i}^{2}}\right)=\sum_{i,j}\left(a_{i,j}\left(\frac{1}{d_{i}d_{j}}-\frac{1}{d_{i}^{2}}\right)\right).$
Since $A$ is symmetric we can rewrite this expression as
$\frac{1}{2}\sum_{i,j}\left(a_{i,j}\left(\frac{2}{d_{i}d_{j}}-\frac{1}{d_{i}^{2}}-\frac{1}{d_{j}^{2}}\right)\right)=-\frac{1}{2}\sum_{i,j}\left(a_{i,j}\left(\frac{1}{d_{i}}-\frac{1}{d_{j}}\right)^{2}\right)\leq 0,$
and the result follows.
7 A Global Geometric Mean Friendship Paradox
In this section we will make use of the following lemma.
Lemma 7.1.
Given a strictly increasing function $f:\mathbb{R}\to\mathbb{R}$, the generalized global friendship inequality (3) holds for $x_{i}=f(d_{i})$, with equality if and only if the network is regular.
Proof.
Since $f$ is strictly increasing, we have, for any $i$ and $j$, $\left(d_{i}-d_{j}\right)\left(f(d_{i})-f(d_{j})\right)\geq 0$, with equality if and only if $d_{i}=d_{j}$. Summing, we obtain,
$\sum_{i}\sum_{j}\left(d_{i}-d_{j}\right)\left(f(d_{i})-f(d_{j})\right)\geq 0,$
with equality if and only if the network is regular. Upon simplification and rearrangement, this inequality reduces to
$\sum_{i}\frac{d_{i}f(d_{i})}{\mathbf{1}^{T}\mathbf{d}}-\frac{1}{n}\sum_{i}f(d_{i})\geq 0,$
as required. ∎
As we mentioned in section 2, Eom and Jo [14] showed that (3) is equivalent to asking for $\mathbf{x}$ to be nonnegatively correlated with the degree vector. Lemma 7.1 shows the intuitively reasonable result that applying an increasing function to the degree vector respects the inequality.
The aim of this section is to consider using the geometric mean in (2) rather than the arithmetic mean. As mentioned in the start of subsection 2.1, the friend-of-friend list includes each degree value $d_{i}$ exactly $d_{i}$ times, and the list involves $\mathbf{1}^{T}\mathbf{d}$ numbers. So the friend-of-friend geometric mean is
$\left(\prod_{i=1}^{n}{d_{i}}^{d_{i}}\right)^{1/\mathbf{1}^{T}\mathbf{d}}.$
The friend geometric mean is
$\left(\prod_{i=1}^{n}{d_{i}}\right)^{1/n}.$
So we may define the global geometric mean friendship inequality to be
$\left(\prod_{i=1}^{n}{d_{i}}^{d_{i}}\right)^{1/\mathbf{1}^{T}\mathbf{d}}\geq\left(\prod_{i=1}^{n}{d_{i}}\right)^{1/n}.$ (22)
Theorem 7.1 (Global geometric mean friendship paradox).
The global geometric mean friendship inequality (22) holds, with equality if and only if the network is regular.
Proof.
Taking logarithms in (22) gives
$\frac{1}{\mathbf{1}^{T}\mathbf{d}}\sum_{i}\log{d_{i}}^{d_{i}}\geq\frac{1}{n}\sum_{i}\log{d_{i}},$
which is equivalent to
$\sum_{i}\frac{d_{i}\log{d_{i}}}{\mathbf{1}^{T}\mathbf{d}}\geq\frac{1}{n}\sum_{i}\log{d_{i}}.$ (23)
The result now follows from Lemma 7.1.
Acknowledgement
DJH was supported by a Leverhulme Fellowship. FH was supported by a Mathematical Institute Scholarship by Oxford Mathematical Institute. FT was partially funded by the PRIN-MUR project MOLE code 2022ZK5ME7 and by PRIN-PNRR project FIN4GEO CUP P2022BNB97.